OpenStudy (anonymous):

which equation can be used to solve the problem?Maura spent 1 and 1/2 hr on science homework.This was 2/3 of the total time she spent on homework.How much time did she spend on homework

OpenStudy (anonymous):

A.h=2/3*1 and 1/2 B.1 and 1/2*h=2/3 C.h-2/3=1 and 1/2 D.2/3h=1 and 1/2

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Here's what I am thinking: 1 1/2 = 2/3 * h where h = total homework time

OpenStudy (anonymous):

okay

Directrix (directrix):

@doopy Does that make sense?

OpenStudy (anonymous):

little

Directrix (directrix):

I see it as the science time is 2/3 the time spent on homework. I translated the words to symbols. 1 1/2 = 2/3 * h where h = total homework time

Directrix (directrix):

I see that in the answer options. Do you?

OpenStudy (anonymous):

yes

Directrix (directrix):

Alrighty, then.